OpenStudy (anonymous):

What is the sum of the geometric series ... A.15,658 B.6,138 C.12,276 D.756

OpenStudy (anonymous):

\[\sum_{n = 1}^{10} 6(2)^x\]

OpenStudy (anonymous):

@Hero Help Please ?

OpenStudy (anonymous):

@Hero I Didnt Block You , But Ok Thanks !

OpenStudy (anonymous):

@satellite73 Help ?

OpenStudy (anonymous):

@JamesJ Help

OpenStudy (anonymous):

@AccessDenied Help

OpenStudy (jamesj):

What is the sum in general of a geometric series? a + ar + ar^2 + ... + ar^(n-1) = ...

OpenStudy (anonymous):

ar ? what does that variable mean

OpenStudy (jamesj):

a and r are variables

OpenStudy (anonymous):

I know , representing what ?

OpenStudy (jamesj):

numbers in a general geometric series. I'm asking you what is the formula for the sum of geometric series so you can use it in this problem. To cut to the chase \[ a + ar + ar^2 + .... + ar^n = \sum_{j=0}^n ar^n = a \frac{r^{n+1} - 1}{r - 1} \] Now apply this in your problem. Be careful about what is a and r in your problem and/or the indices

OpenStudy (jamesj):

\[ a + ar + ar^2 + .... + ar^n = \sum_{j=0}^n ar^j = a \frac{r^{n+1} - 1}{r - 1} \]